# Diversity-Oriented Data Augmentation with Large Language Models

### v1

@@ 4.2 Overall Results
To verify the effectiveness of DoAug, we evaluate downstream task accuracy alongside the diversity and affinity of the augmented dataset. We report the rankings of performance, diversity, and affinity averaged on 12 datasets achieved by DoAug and 11 baseline methods in Figure 2. From these results, we have the following observations: (1) DoAug achieves the highest performance on downstream tasks compared to other SOTA data augmentation methods, as indicated by the color bar in Figure 2. This demonstrates the high quality and superior adaptability of the datasets generated by our proposed method in real-world applications. (2) DoAug achieves the highest diversity score and outperforms all other baseline methods. This implies that DoAug effectively improves dataset diversity. (3) DoAug achieves a considerably high position on the affinity rankings, indicating that the sample semantics are preserved to the greatest extent possible. In sum, DoAug achieves the top position in the combined dataset diversity and affinity rankings. Additionally, it achieves the best downstream task performance, indicated by the lightest yellow.
Distance $\uparrow$ Dispersion $\uparrow$ Radius $\uparrow$ Homogeneity $\uparrow$ Vocabulary $\uparrow$ Average Rankings $\downarrow$
Original 8.9458 (11) .1911 (13) .5352 (6) .8834 (11) 4737 (8) 9.8
OCR 8.8516 (13) .1963 (9) .5198 (13) .8947 (9) 5764 (2) 9.2
Keyboard 8.9439 (12) .1978 (7) .5233 (12) .8990 (8) 5939 (1) 8.0
EDA 9.1927 (6) .2090 (3) .5247 (11) .9010 (7) 4416 (11) 7.6
BT 9.2761 (4) .2050 (4) .5345 (7) .8744 (12) 4782 (7) 6.8
Unmask 9.0021 (10) .1932 (12) .5344 (8) .9041 (6) 4158 (13) 9.8
AugGPT 9.1522 (8) .2005 (6) .5340 (9) .9094 (3) 4585 (10) 7.2
Grammar 9.5250 (2) .2121 (2) .5384 (3) .8332 (13) 4316 (12) 6.4
Spelling 9.0346 (9) .1951 (11) .5315 (10) .9080 (4) 4917 (5) 7.8
Chain 9.1946 (5) .1964 (8) .5385 (2) .9099 (2) 4974 (4) 4.2
Hint 9.3608 (3) .2037 (5) .5388 (1) .8946 (10) 4796 (6) 5.0
Taboo 9.1566 (7) .1953 (10) .5379 (4) .9148 (1) 4690 (9) 6.2
DoAug 9.5308 (1) .2154 (1) .5355 (5) .9054 (5) 5321 (3) 3.0
Table 2: 5 Diversity metrics averaged on 12 datasets.
@@ 4.3.1 Performance Gains
The full results for BERT classification performance on original and augmented datasets are presented in Table 1. The results show that DoAug surpasses all baseline methods on average. Specifically, it outperforms the baseline methods on 11 out of 12 datasets except on RTE. DoAug achieves performance gain of $10.52\%$ on average, surpassing the runner-up method, namely Taboo, with 3.76 percentage points.
@@ 4.3.2 Diversity Gain
We demonstrate the diversity gains in terms of all 5 diversity metrics along with their rankings achieved by DoAug and baseline methods in Figure 2. DoAug ranks the top on the chart with an average ranking of 3.0. Specifically, it achieves the best for the Distance and Dispersion metrics. For the Vocabulary metric, OCR and Keyboard include large amounts of out-of-vocabulary tokens due to character level perturbation. The three baselines with diversity incentives, namely Chian, Hint, and Taboo, also achieve reasonably good diversity gain, in line with the results of Cegin et al. (2024).
@@ 4.3.3 Affinity and Paraphrase Validity
We present the affinity of DoAug and baseline methods in Figure 3, where DoAug outperforms other methods except Unmask, whose affinity score is 1.95. The Unmask method generates augmentation by replacing randomly selected words with “[MASK]” and predicts the masked words with the BERT model. Since the augmented samples are recovered from the BERT embeddings of the corrupted original samples and we use the BERT embeddings to calculate affinity, it is reasonable to yield extremely high affinity scores. Following Cegin et al. (2023, 2024), we also investigate paraphrase validity at the sample level. We prompt the DeepSeek-V3 model to check if the paraphrases are semantically similar to the original samples and adhere to the original labels. Results show that $97\%$ paraphrases are valid, suggesting the data augmenter introduces negligible noises to the dataset.
@@ 4.4 Ablation Studies
To verify the effectiveness of our proposed method, we conduct ablation studies to show that all components in the method framework contribute to the final performance gains, as shown in Table 3, where w/o Coreset refers to applying augmentation on a random subset of the dataset without deriving a coreset of importance samples, w/o Selective refers to augmenting samples in both $\mathcal{S}_{\text{retain}}$ and $\mathcal{S}_{\text{augment}}$ instead of only augmenting the latter, w/o Aug refers to using the coreset directly for training without data augmentation, w/o DPO refers to using the LLM paraphraser from the SFT stage for data augmentation, w/o DS refers to removing the diversity-based sampling module, and w/o DPO|DS removes both steps. Thus, all components in our method framework contribute to the final performance gains. As Figure 4 shows, we also study the effect of diversity-oriented fine-tuning and diversity-based sampling on the dataset diversity, demonstrating that all proposed components are effective.
@@ 4.5 LLM Architectures Exploration
For the LLM augmenter, we replace the Llama-3.2-1B-Instruct model with the similar-sized Qwen2.5-1.5B-Instruct model Team (2024). As shown in Figure 5, datasets augmented by the Qwen model significantly outperform the original dataset and achieve comparable accuracy and diversity with those of the Llama model. Detailed results are given in Appendix F.
@@ 5 Conclusion
Diversity is an important factor in developing AI-ready, high-quality datasets but is often ignored in data augmentation methods. We propose a Diversity-oriented data Augmentation framework (DoAug) that trains an LLM paraphraser to enlarge and introduce diversity to textual datasets. The LLM paraphraser is fine-tuned to rewrite existing sentences in the original datasets, generating high-affinity samples and ensuring coherence of the dataset. We further construct a preference dataset and then fine-tune the LLM paraphraser with the DPO algorithm to encourage diversified generation. In this way, we maximize the diversity of the augmented dataset in our method. In extensive experiments, our proposed method exhibits a remarkable capability to boost dataset diversity, and the diversity gain significantly benefits the model’s learning performance of downstream tasks.
@@ 6 Limitations
Diversity Exploration: The evaluation of diversity lacks a universally accepted metric. In this study, we employed a subset of diversity-related evaluation methods, but other metrics, such as n-gram-based measures (e.g., assessing uniqueness or non-repetition), were not utilized. This limitation suggests that our assessment of diversity may not fully capture all aspects of the concept.
Augmentation Validation: Evaluating the correctness of generated data remains a challenging task. While both human evaluation and model-assisted evaluation are viable approaches, each comes with its own limitations. In this study, we relied on a sentiment-neutral large language model (LLM) for evaluation. However, this approach has inherent constraints, such as potential biases in the model and its inability to fully capture nuanced correctness in certain contexts.
Generation Factors: The quality and characteristics of generated samples are influenced by multiple factors, including the choice of prompts and the specific LLMs used. In this study, the prompts and LLMs were limited in number, and we did not exhaustively explore all possible configurations. This restriction may have impacted the diversity and quality of the generated samples.
Evaluation Benchmarks: Our evaluation was primarily focused on sentence classification tasks, and we did not extend our analysis to more general tasks, such as mathematical reasoning, instruction-following, creative writing, question-answering (QA), or chain-of-thought (CoT) reasoning. Additionally, we did not explore multimodality scenarios, which could provide a broader perspective on the applicability of our framework. Furthermore, the datasets used in this study were limited to English corpora, and we did not consider multilingual datasets, which could offer insights into cross-lingual or language-specific performance.
@@
These limitations highlight areas for future work, including the adoption of more comprehensive diversity metrics, broader validation approaches, exploration of additional prompts and LLMs, and evaluation across diverse tasks, modalities, and languages.
Appendix A Dataset Specification
The details of the 12 NLP datasets used in our experiments, including the domain, application task, input scheme, and class number, are summarized in Table 5.
Appendix B Implementation Details
@@
We use Llama-3.2-1B-Instruct with BF16 quantization as the LLM paraphraser. Llama-3.2 is one of the latest products of the Llama family. The Llama-3.2-1B model outperforms the Llama-3.2-3B and Llama-3.1-8B models on the rewriting task while requiring minimal memory and inference time [^2] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct. We fine-tune the Llama model with the LlamaFactory framework Zheng et al. (2024). In the SFT stage, the learning rate is $10^{-4}$. In the DPO stage, $\beta$ in the loss function is set to $0.1$ and the learning rate is $5^{-6}$. The LLM is trained for 3 epochs with the AdamW optimizer, a cosine scheduler, and a warm-up ratio of $0.1$ in each stage. The rank $r$ for LoRA fine-tuning is $8$. $\mathcal{D}_{\text{SFT}}$ and $\mathcal{D}_{\text{DPO}}$ contain 100,000 sentence pairs and 50,000 preference pairs, respectively. We use the embedding vector of the [CLS] token in the last layer of the $\text{BERT}_{\text{base}}$ model as the embedding space $\mathcal{E}$ for both preference dataset construction and dataset diversity evaluation. In the coreset selection step, the coreset ratio $r_{\text{augment}}:r_{\text{retain}}:r_{\text{prune}}$ is $1:1:1$. For each original sentence, we generate $K=5$ paraphrases and sample the most diversified output. For downstream task evaluation, we train the BERT model for 3 epochs. The model is updated with the AdamW optimizer. The learning rate is $5^{-5}$ with a linear scheduler and no warm-up. The LLM paraphraser and downstream task models are trained on two A100-40G GPUs with transformers 4.45.2, pytorch 2.5.1, and CUDA 12.4. It takes 32 and 36 minutes to train with SFT and DPO, respectively, and the LLM augmenter can paraphrase roughly 1 sentence per second.
Appendix C Baseline Methods Implementation
@@
The error ratios for OCR and Keyboard are set to $0.15$. The ratios for its EDA’s four operations in our experiments are set to $0.1$, respectively. For BT, we use the en-zh and zh-en versions of the opus-mt model by Helsinki-NLP Tiedemann and Thottingal (2020) as the translator. The masking ratio for Unmask is set to $0.15$ and we sample the top-1 predictions of the $\text{BERT}_{\text{base}}$ model. Considering the cost of calling ChatGPT APIs, when implementing AugGPT, Grammar, Spelling, Chain, Hint, and Taboo, we use the open-source Llama-3.1-8B-Instruct model Dubey et al. (2024) as a replacement. Since it achieves competitive performance compared with the GPT 3.5 Turbo model (e.g. 69.4 v.s. 70.7 on MMLU 5-shot and 80.4 v.s. 69.9 on IFEval) Dubey et al. (2024), replacing GPT with Llama hardly compromises the effectiveness of the baseline methods.
Appendix D Augmentation Examples

### v2

@@ 4.2 Overall Results
To verify the effectiveness of DoAug, we evaluate downstream task accuracy alongside the diversity and affinity of the augmented dataset. We report the rankings of performance, diversity, and affinity averaged on 12 datasets achieved by DoAug and 12 baseline methods in Figure 3. From these results, we have the following observations: (1) DoAug achieves the highest performance on downstream tasks compared to other SOTA data augmentation methods, as indicated by the color bar in Figure 3. This demonstrates the high quality and superior adaptability of the datasets generated by our proposed method in real-world applications. (2) DoAug achieves the highest diversity score and outperforms all other baseline methods. This implies that DoAug effectively improves dataset diversity. (3) DoAug achieves a considerably high position on the affinity rankings, indicating that the sample semantics are preserved to the greatest extent possible. In sum, DoAug achieves the top position in the combined dataset diversity and affinity rankings. Additionally, it achieves the best downstream task performance, indicated by the lightest yellow.
Figure 3: Diversity, affinity, and performance achieved by DoAug and baseline methods. Results are averaged on 12 datasets and the diversity rankings are further averaged on 6 metrics in this diagram. A smaller number for the rankings indicates better results.
@@ 4.3.1 Performance Gains
The full results for BERT classification performance on original and augmented datasets are presented in Table 1. The results show that DoAug surpasses all baseline methods on average. Specifically, it outperforms the baseline methods on 11 out of 12 datasets except on RTE. DoAug achieves performance gain of $10.52\%$ on average, surpassing the runner-up method with an advantage of 3.76 percentage points.
@@ 4.3.2 Diversity Gain
We demonstrate the diversity gains in terms of all 6 diversity metrics achieved by DoAug and baseline methods in Figure 2. For each metric, the scores are normalized to $[0,1]$, and we also include the original scores in Appendix F. DoAug ranks the top on the chart with an average score of 0.98. Specifically, it achieves the best for the Distance, Dispersion, Vocabulary Size, and Unique 3-grams metrics. It is also competitive in terms of Isocontour Radius and Homogeneity. The three baselines with diversity incentives, namely Chain, Hint, and Taboo, also achieve reasonably good diversity gain, in line with the results of Cegin et al. (2024).
@@ 4.3.3 Affinity and Paraphrase Validity
We present the affinity of DoAug and baseline methods in Figure 4, where DoAug outperforms other methods except Unmask, whose affinity score is 1.95. The Unmask method generates augmentation by replacing randomly selected words with “[MASK]” and predicts the masked words with the BERT model. Since the augmented samples are recovered from the BERT embeddings of the corrupted original samples and we use the BERT embeddings to calculate affinity, it is reasonable to yield extremely high affinity scores. Following Cegin et al. (2023, 2024), we also investigate paraphrase validity at the sample level. We perform a human evaluation on 200 samples per dataset to check if the paraphrases are semantically similar to the original samples and adhere to the original labels. Results show $95\%$ paraphrases are valid. We also prompt the DeepSeek-V3 model, a very strong and knowledgeable LLM that is good at understanding users’ intent and evaluating the task scenario for an LLM-based evaluation. Results show that $97\%$ paraphrases are valid, suggesting that DoAug introduces negligible noises to the dataset.
@@ 4.4 Ablation Studies
To verify the effectiveness of our proposed method, we conduct ablation studies as shown in Table 3, where w/o Coreset refers to applying augmentation on a random subset of the dataset without deriving a coreset of importance samples, w/o Selective refers to augmenting samples in both $\mathcal{S}_{\text{retain}}$ and $\mathcal{S}_{\text{augment}}$ instead of only augmenting the latter, w/o Aug refers to using the coreset directly for training without data augmentation, w/o DPO refers to using the LLM paraphraser from the SFT stage for data augmentation, w/o DS refers to removing the diversity-based sampling module, and w/o DPO|DS removes both steps. From these results, we find out that all components in the method framework make remarkable contributions to the final performance gains. We also notice that focusing augmentation on the selected coreset is the most important factor of performance (with the lowest average scores and no runner-up results). As Figure 5 shows, we also study the effect of diversity-oriented fine-tuning and diversity-based sampling on the dataset diversity, demonstrating that all proposed components are effective. Concretely, we can observe that diversity sampling has more influence on sample-level latent diversity, while DPO training has more influence on lexical diversity. Further, we investigate whether the DPO training can be replaced by cost-saving approaches, such as sampling from a higher temperature and using prompts with diversity incentives. Results in Figure 6 indicate that replacing DPO training with sampling from a higher temperature or using prompts with diversity incentives achieves inferior results, failing to compete with the DPO version, suggesting model fine-tuning is necessary for diversity gain and performance improvement. Full results are in Appendix H.
@@
(a) Accuracy
(b) Distance
(c) 3-grams
Figure 6: Replacement study on DPO training.
@@ 4.5 LLM Architectures Adaptability
For the LLM augmenter, we replace the Llama-3.2-1B-Instruct model with the similar-sized Qwen2.5-1.5B-Instruct model Team (2024). As shown in Figure 7, datasets augmented by the Qwen model significantly outperform the original datasets in diversity and achieve comparable accuracy with those of the Llama model. Detailed results are given in Appendix L.
@@ 5 Conclusion
Diversity is an important factor in developing AI-ready and high-quality datasets but is often ignored in data augmentation methods. We propose an innovative Diversity-oriented data Augmentation framework (DoAug) that fine-tunes an LLM paraphraser to enlarge and introduce diversity to textual datasets. The LLM paraphraser is fine-tuned to rewrite existing seed sentences in the original datasets, generating high-affinity samples and ensuring coherence of the dataset. We further construct a preference dataset and then fine-tune the LLM paraphraser with the DPO algorithm to encourage diversified generation. In this way, we maximize the diversity of the augmented dataset in our method. In extensive experiments, our proposed method exhibits a remarkable capability to boost dataset diversity, and the diversity gain significantly benefits the model’s learning performance of downstream tasks.
6 Acknowledgments
Pengfei Wang is supported by the National Natural Science Foundation of China (Grant Nos. 62406306 and 92470204), and the Science and Technology Development Fund (FDCT), Macau SAR (file no. 0123/2023/RIA2, 001/2024/SKL).
@@ 7 Limitations
Diversity Exploration: The evaluation of diversity lacks agreement on universally accepted metrics. In this study, we employed a subset of diversity-related evaluation methods, but other metrics, such as human-centered diversity evaluation, were not utilized. This limitation suggests that our assessment of diversity may not fully capture all aspects of the concept.
Augmentation Validation: Evaluating the correctness of generated data remains a challenging task. While both human evaluation and model-assisted evaluation are viable approaches, each comes with its own limitations. In this study, we employed both a task-aware LLM and humans for evaluation. However, there is no perfect human or LLM and this approach has inherent constraints, such as potential biases in the LLMs’ training corpus or humans’ knowledge and their inability to fully capture nuanced correctness in certain contexts.
Generation Factors: The quality and characteristics of generated samples are influenced by multiple factors, including the generation temperature, the choice of prompts, and the specific LLMs used. In this study, for each above-mentioned factor, we only explored two different settings, and we did not exhaustively explore all possible configurations. This restriction may have impacted the diversity and quality of the generated samples.
Evaluation Benchmarks: Our evaluation was primarily focused on sentence classification tasks and included two QA-based reasoning tasks, and we did not extend our analysis to more general tasks, such as mathematical reasoning, instruction-following, creative writing, or chain-of-thought (CoT) reasoning. Additionally, the datasets used in this study concentrated on English corpora, and we only considered one multilingual dataset, which could offer insights into cross-lingual or language-specific performance. Furthermore, we did not explore multimodality scenarios, which could provide a broader perspective on the applicability of our framework.
@@
These limitations highlight potential areas for future work, especially the adoption of more comprehensive diversity metrics and evaluation across diverse data modalities.
Appendix A Evaluation Criterion
$\bullet$ Distance assesses the average distance between samples as follows: ${Distance(\mathcal{S})=\frac{1}{|\mathcal{S}|}\sum_{x_{i},x_{j}\in\mathcal{S}}\sqrt{(e_{x_{i}}-e_{x_{j}})^{2}}}$, where $e_{x}=\mathcal{E}(x)$ is the embedding of sample $x$ in the embedding space $\mathcal{E}$, and a larger distance indicates greater diversity.
[…]
The Distance, Dispersion, Isocontour Radius, and Homogeneity scores are calculated class-wise and then averaged over all classes, while vocabulary size and Unique 3-grams are directly derived from the entire dataset. Invalid (wrong) words are excluded when calculating lexical diversity.
[…]
Domain Application task Input #Classes
ANLI Nie et al. (2020) General Entailment annotation Sentence pair 3
ChemProt Kringelum et al. (2016) Medical Chemical-protein relationship Single sentence 13
CoLA Wang et al. (2019) General Acceptability judgment Single sentence 2
MNLI Wang et al. (2019) General Entailment annotation Sentence pair 3
MPQA Wiebe et al. (2005) General Sentiment analysis Single sentence 2
MRPC Wang et al. (2019) General Semantically equivalence Sentence pair 2
RCT Dernoncourt and Lee (2017) Medical Role of sentence Single sentence 5
RTE Wang et al. (2019) General Entailment annotation Sentence pair 2
SST-2 Wang et al. (2019) General Sentiment analysis Single sentence 2
SUBJ Pang and Lee (2004) General Subjective v.s. objective Single sentence 2
Symptoms Dai et al. (2025) Medical Disease judgment Single 25
Yelp Zhang et al. (2015) General Review rating Single 5
Table 5: A summary of 12 textual datasets.
Symptoms SST-2
Original My joints ache whenever it is cold I have always appreciated a smartly written motion picture
OCR My joint5 aehe when€ver it is co1d I hawo alwa9\$ appraciateol o smartly wri7tcn mo7ion pic+ure
Keyboard My joibts axhe wjrnever it js cild I have alwats apprecuated a smartly writren motion picthre
EDA My joints ache whenever it is common cold I have always liked a smartly written motion picture
AEDA My joints : ache whenever , it is ? cold I have . always appreciated ; a smartly written . motion picture
BT My joint hurts when it’s cold I always admire smart writing action pictures
Unmask My joints ache. it is cold I have always appreciated a smartly written motion picture
AugGPT I experience joint pain during cold weather I’ve always been fond of motion pictures that showcase clever writing
Grammar Whenever it is cold, my joints ache A motion picture that has been always appreciated by me is smartly written
Spell My joints ake whenever it is colld I have allwas apreciated a smartly written motion picture
Chain I have trouble with my joints in cold weather I am a fan of movies that are skillfully made and have a captivating storyline
Hint Cold weather often makes my joints feel stiff I have a fondness for motion pictures that are well-written, well-crafted, and have a long-standing appreciation for them
Taboo Whenever it gets chilly, my joints feel quite sore I have a fondness for motion pictures that are well-crafted
DoAug I have trouble moving my joints in cold weather, causing discomfort I have a fondness for films that are well-crafted and have a sophisticated style
Table 6: Augmentation examples on the Symptoms and SST-2 datasets.
Appendix B Dataset Specification
The details of the 12 NLP datasets used in our experiments, including the domain, application task, input scheme, and class number, are summarized in Table 5. ANLI Nie et al. (2020), MNLI Wang et al. (2019), and RTE Wang et al. (2019) are three datasets that require models to recognize the textual entailment of two sentences, which requires decent reasoning ability of models. ChemProt Kringelum et al. (2016), RCT Dernoncourt and Lee (2017), and Symptoms Dai et al. (2025) are three medical datasets that involve domain knowledge of the models. ChemProt describes the relationship between chemical-protein Paris, RCT requires the model to analyze what role a sentence plays in the abstract of a medical research paper, and the Symptoms dataset is about judging the disease from patient complaints. MPQA Wiebe et al. (2005), SST-2 Wang et al. (2019), and Yelp Zhang et al. (2015) and sentiment datasets, where MPQA and SST-2 label the sentences as “negative” or “positive”, and Yelp assigns numerical ratings from 1 to 5 to the reviews. CoLA Wang et al. (2019) evaluates the acceptability of a sentence, MRPC Wang et al. (2019) evaluates if a pair of sentences are equivalent, and SUBJ Pang and Lee (2004) evaluates if a sentence is subjective or objective. Symptoms dataset is available at Kaggle [^3] https://www.kaggle.com/datasets/paultimothymooney/medical-speech-transcription-and-intent. GLUE benchmark datasets (CoLA, MNLI, MRPC, RTE, and SST-2) [^4] https://huggingface.co/datasets/nyu-mll/glue, ANLI44footnotetext: https://huggingface.co/datasets/facebook/anli, ChemProt [^5] https://huggingface.co/datasets/AdaptLLM/ChemProt, RCT [^6] https://huggingface.co/datasets/AdaptLLM/RCT, MPQA [^7] https://huggingface.co/datasets/rahulsikder223/SentEval-MPQA, SUBJ [^8] https://huggingface.co/datasets/SetFit/subj, and Yelp [^9] https://huggingface.co/datasets/Yelp/yelp_review_full are also available at Hugging Face.
Appendix C Implementation Details
@@
We use Llama-3.2-1B-Instruct with BF16 quantization as the LLM paraphraser. Llama-3.2 is one of the latest products of the Llama family. The Llama-3.2-1B model outperforms the Llama-3.2-3B and Llama-3.1-8B models on the rewriting task while requiring minimal memory and inference time [^10] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct. The prompt for paraphrasing is “You will be given a sentence. Please paraphrase the sentence.” We fine-tune the Llama model with the LlamaFactory framework Zheng et al. (2024). When training the model, we use LoRA to reduce computation costs. We train a LoRA adapter for the SFT stage, merge the SFT adapter, train a LoRA adapter for the DPO stage, and finally merge the DPO adapter for use. In the SFT stage, the learning rate is $10^{-4}$. In the DPO stage, $\beta$ in the loss function is set to $0.1$ and the learning rate is $5^{-6}$. The LLM is trained for 3 epochs with the AdamW optimizer, a cosine scheduler, and a warm-up ratio of $0.1$ in each stage. The rank $r$ for LoRA fine-tuning is $8$. $\mathcal{D}_{\text{SFT}}$ contains 100,000 sentence pairs (20,000 original sentences and 5 paraphrases for each of them). In the SFT stage, the model is trained to produce one paraphrase for one input. $\mathcal{D}_{\text{DPO}}$ contains 50,000 preference pairs. We use the embedding vector of the [CLS] token in the last layer of the $\text{BERT}_{\text{base}}$ model as the embedding space $\mathcal{E}$ for both preference dataset construction and dataset diversity evaluation. In the coreset selection step, the coreset ratio $r_{\text{augment}}:r_{\text{retain}}:r_{\text{prune}}$ is $1:1:1$. For each original sentence, we generate $K=5$ paraphrases and sample the most diversified output. For downstream task evaluation, we train the BERT model for 3 epochs. The model is updated with the AdamW optimizer. The learning rate is $5^{-5}$ with a linear scheduler and no warm-up. Experiments are repeated with ten random seeds. The LLM paraphraser and downstream task models are trained on two A100-40G GPUs with transformers 4.45.2, pytorch 2.5.1, and CUDA 12.4. Training with SFT and DPO takes 32 and 36 minutes, respectively, and the LLM augmenter can paraphrase roughly one sentence per second.
Appendix D Baseline Methods Implementation
@@
The error ratios for OCR and Keyboard are set to $0.15$. The ratios for EDA’s four operations are set to $0.1$. AEDA’s punctuation ratio is $0.3$ as used in the original method. For BT, we use the en-zh and zh-en versions of the opus-mt model by Helsinki-NLP Tiedemann and Thottingal (2020) as the translator. The masking ratio for Unmask is set to $0.15$ and we sample the top-1 predictions of the $\text{BERT}_{\text{base}}$ model. Considering the cost of calling ChatGPT APIs, when implementing AugGPT, Grammar, Spelling, Chain, Hint, and Taboo, we use the open-source Llama-3.1-8B-Instruct model Dubey et al. (2024) as a replacement. Since it achieves competitive performance compared with the GPT 3.5 Turbo model (e.g. 69.4 v.s. 70.7 on MMLU 5-shot and 80.4 v.s. 69.9 on IFEval) Dubey et al. (2024), replacing GPT with Llama hardly compromises the effectiveness of the baseline methods. For the original dataset and all baselines, we also set the coreset ratio $r_{\text{augment}}:r_{\text{retain}}:r_{\text{prune}}$ to $1:1:1$ but do not rank the samples before selecting; for the original dataset, we do not augment $\mathcal{S}_{\text{augment}}$. In this way, the number of samples from the original dataset is the same for all methods, ensuring fairness in terms of the samples’ coverage and distribution. The number of final samples is in accordance with DoAug for all baselines, further ensuring the fairness of model training and evaluation. We use the same random seeds for all baselines as for ours.
Appendix E Augmentation Examples